OpenStudy (anonymous):

1 meter3= centimeters3

OpenStudy (kira_yamato):

1 m3 = (100)^3 cm3 = 10^6 cm3

OpenStudy (anonymous):

1 m = 100 cm So, 1 cubic m = 1000000 cubic cm

OpenStudy (anonymous):

(100*100*100) = 1,000,000 cm ^3

OpenStudy (anonymous):

correct